# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v2

@@ II-B Uplink Transmission and Spectral Efficiency
Let $x_{t}$ denote the unit power uplink data signal transmitted by UE $t$. The uplink normalised data power corresponding to UE $t$ and additive white Gaussian noise are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively. The uplink signal received at the $m$-th AP is:
$\textbf{y}^{u}_{m}=\sum\nolimits_{t=1}^{T}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
In decentralized data detection, unlike centralized approach where each AP sends received signal and channel estimates to CPU, each AP performs data detection for its served UEs using a local linear received combining vector. Following the proposed decentralized decoding architecture, each AP computes its own interference-suppressing local weight based on the local signal-to-interference-plus-noise ratio (SINR). These weights are applied to the locally detected data symbols before being sent to the CPU. This differs from LSFD, where local data and channel estimates are transmitted to the CPU for centralized processing. The locally weighted data symbol $\hat{x}_{mt}$ for UE $t$ at AP $m$ is given by
$\hat{x}_{mt}=a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m},$ (2)
where $\mathbf{v}_{m{i_{t}}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for UE $t$ at AP $m$, and $a_{mt}$ is the corresponding local weight. These weighted estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
$\hat{x}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m}.$ (3)
After expanding $\textbf{y}^{u}_{m}$, the (3) can also be rewritten as
$\begin{split}\hat{x}_{t}&=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}+\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\tilde{\textbf{g}}_{mt}\\ &+\sum_{\begin{subarray}{c}k=1\\ k\neq t\end{subarray}}^{T}\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}x_{k}+\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$ (4)
Here, the first term is the desired signal for UE $t$, the second term is interference term due to imperfect channel estimation, the third term is interference term due to all other UEs and last term is the noise. Based on the decomposition in (4), the achievable uplink spectral efficiency is obtained using the bounding technique and the corresponding Shannon capacity lower bound [17, 18, 16] is given by Theorem 1.
@@
where,
$\begin{split}&\textsc{DS}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}\right\},\\ &\textsc{UI}_{tk}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk},\\ &\textsc{GN}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$
@@
III Adaptive Combining and Local Weights Analysis
In this section, we first derive closed-form expressions for the uplink SE and local combining weights for both G-PFZF and G-PWPFZF schemes. We also formulate an optimization problem to partition the pilots into strong and weak.
@@ III Adaptive Combining and Local Weights Analysis
III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
In [13], the combining vectors for UE $t$ is dependent on $\beta_{mt}$, which is unique to each UE. As a result, generating a distinct combining vector for every UE entails a number of multiplications that scales linearly with the total number of UEs. To reduce computational cost, we propose G-PFZF combining scheme, where each AP independently designs a combining vector for each of its pilots and pilot-sharing UEs use the same pilot, as APs are not able to distinguish among individual UEs sharing the same pilot. The design of combining vectors depends on whether a pilot is strong or weak. Thus, each AP categorizes the pilots into two groups, strong and weak based on the local sum SE optimization. Unlike traditional PFZF, G-PFZF does not force the APs to select atleast one pilot as strong and allows the APs to choose the combining schemes among PFZF or MR or FZF.
Remark 2: In our proposed combining schemes, a strong pilot indicates that a ZF combining vector is assigned to that pilot (and its corresponding UEs), while a weak pilot is assigned an MR combining vector.
For AP $m$, a set of strong pilots is denoted by $\mathcal{S}_{m}$ and a complementary set of weak UEs as $\mathcal{W}_{m}$. Let $L_{\mathcal{S}_{m}}$ denote the number of distinct strong pilots used in set $\mathcal{S}_{m}$, and define the corresponding set of pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}{=}\{r_{m,1},r_{m,2},{\dots},r_{m,L_{\mathcal{S}_{m}}}\}$. The matrix $\mathbf{\bar{G}}_{m}{\in}\mathbb{C}^{A\times L_{p}}$ contains the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilot sequences, whereas the strong pilots utilize only $L_{\mathcal{S}_{m}}$ of them. To isolate the relevant columns of $\mathbf{\bar{G}}_{m}$, we define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}{=}[\mathbf{e}_{r_{m,1}},{\dots},\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]{\in}\mathbb{C}^{L_{p}\times L_{\mathcal{S}_{m}}}$, where $\mathbf{e}_{r}$ denotes the $r$-th column of the identity matrix $\mathbf{I}_{L_{p}}$. The resulting product $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$ yields a full-rank matrix containing only $L_{\mathcal{S}_{m}}$ linearly independent columns corresponding to the strong pilots. For a given pilot $i{\in}\mathcal{S}_{m}$, let $j_{mi}{\in}\{1,2,\dots,L_{\mathcal{S}_{m}}\}$ denote the index of pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}{\in}\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}{=}\mathbf{e}_{i}$.
@@ III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
Following the zero-forcing principle, the pseudo-inverse of this matrix can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$. To obtain the combining vector associated with a specific pilot, this pseudo-inverse term is multiplied by $\bm{\varepsilon}_{j_{mi}}$ and subsequently normalized. Thus, the local ZF combining vector for pilot $i{\in}\mathcal{S}_{m}$ at AP $m$ is given by:
@@
Following [19, 20, 21], the normalization term in (7), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix ($\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$) with degree of freedom satisfying $A\geq L_{\mathcal{S}_{m}}+1$, can be rewritten as
@@
$\begin{split}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (11)
@@
Note 1: From (11) and (13), it is evident that local ZF achieves superior interference suppression compared to MR, since each AP can eliminate interference among UEs with pilots grouped as strong. However, this interference mitigation comes at the cost of $L_{\mathcal{S}_{m}}$ degrees of freedom, which reduces the array gain available for desired signal enhancement.
By using the combining vectors in (7) and (8), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by
@@
The derivation of the SINR expression is given in Appendix -A.
Local Weight Design for G-PFZF Scheme: A fundamental limitation of the conventional LSFD architecture is its reliance on centralized optimization [13]. The optimal LSFD weights are calculated at the CPU using knowledge of global channel statistics, which creates a significant fronthaul signaling overhead and computational burden, thus undermining the scalability of the network. In this work, we break this dependency by introducing a novel decentralized decoding architecture where each AP $m$ independently computes a local interference-suppressing weight $a_{mt}$ for each UE $t$, based solely on its local channel information. Since these weights are derived from the LSFCs, they evolve much slowly than the small-scale fading components, ensuring long-term stability and minimal operational overhead.
The primary objectives of this local weight design are twofold: (1) to maximize the contribution of the local estimate to the final SINR at the CPU, and (2) to actively mitigate the dominant interference components, specifically, pilot contamination and inter-user interference, at their source, prior to transmission. This approach eliminates the need for the APs to send the channel estimates to the CPU and the CPU to perform any complex optimization, reducing its role to that of a simple aggregator that sums the pre-weighted estimates from all APs.
The main idea behind designing decoding weights is that it should amplify the data signal received from the APs with strong local desired signal component and suppress those dominated by interference. In LSFD design, weights are computed through coordination among the APs, specifically the coherent interference terms. In contrast, the proposed decentralized framework eliminates inter-AP coordination, requiring each AP to determine its weight independently. Accordingly, an AP with a higher local SINR is given a larger weight than the one with a lower SINR in order to enhance the global SINR. For simplicity, the local weight is chosen to be equal to the local SINR.
The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -A. Accordingly, the closed-form of the local SINR or local weight with respect to AP $m$ of UE $t$ is given by:
$a_{mt}=\text{SINR}^{\text{G-PFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (16)
where
@@
${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
The numerator represents the desired signal strength for the desired UE $t$ at AP $m$, capturing the gain achieved through the chosen combining strategy. The denominator is designed to penalize the weight by two main sources of impairment: the coherent interference from pilot contamination and the combined power of non-coherent interference due to channel estimation error and noise.
By applying these weight to its local estimate, each AP independently produces a soft decision that is already pre-optimized for local interference suppression. This design ensures that the signals forwarded to the CPU require no further large-scale fading decoding, thereby completely eliminating the associated fronthaul overhead for channel statistics and the computational cost of centralized weight calculation. The result is a radically simplified and highly scalable uplink processing architecture that retains the performance benefits of distributed APs design.
Adaptive Pilot Grouping for G-PFZF Combining: This subsection details the local optimization framework that each AP employs to independently partition its set of pilots into strong and weak groups. This classification dictates the appropriate combining strategy, G-PFZF for strong pilots or MR for weak ones, and crucially, generalizes to pure FZF or pure MR should the optimization deem them to be optimal.
The basis of this optimization is the local SINR expression given in (16). As $L_{\mathcal{S}m}{=}\sum_{i=1}^{L_{p}}\delta_{mi}$ and substituting $L_{\mathcal{S}m}$ into (16) makes the dependency explicit as given in (17).
@@
This expression reveals a fundamental trade-off governed by the size of the strong pilot set, $L_{\mathcal{S}m}$. Allocating a pilot to the strong group ( $\delta_{mi}{=}1$) employs ZF to suppress the non-coherent interference from that pilot’s UEs as well as other strong pilots UEs. However, this comes at the cost of a reduced array gain, $(A-L_{\mathcal{S}_{m}})$, for all UEs served by the AP, which diminishes the desired signal power. Conversely, classifying a pilot as weak preserves the full array gain but forgoes any interference suppression for its UEs. This inherent trade-off between maximizing signal strength and minimizing interference makes the grouping problem non-trivial and renders a fixed, network-wide threshold highly suboptimal.
To navigate the above mentioned trade-off, AP $m$ independently solves a local optimization problem, which is formulated as
@@
where $\bm{\delta}_{m}=[\delta_{m1},\delta_{m2},\cdots,\delta_{mL_{p}}]^{\intercal}$. The constraint (18c) represents the necessary condition of Wishart matrix for interference suppression. Also, intuitively, this would otherwise lead to invalid negative values for the desired signal power and coherent interference in the SINR expression. The optimization problem in (18) is a Mixed-Integer Non-Linear Program (MINLP), a class of problems known for its exponential computational cost. To develop a more tractable formulation, we begin by relaxing the binary constraint (18b) into a continuous one:
@@
To ensure that the solutions of this new NLP satisfy the original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the following complementary constraint:
@@
This problem can be solved using non-convex solvers or the successive convex approximation optimization method. Both of these are computationally expensive for large network deployment as non-convex solvers. Therefore, to solve this problem with low computation cost, we employ gradient based proximal gradient ascent (PGA) method [22, 23]. We first reformulate the problem to be unconstrained except for the box constraint (19). The reformulated problem is:

### v3

@@ II-B Uplink Transmission and Spectral Efficiency
Let $x_{t}$ denote the unit power uplink data signal transmitted by UE $t$. The uplink normalised data power corresponding to UE $t$ and AWGN vector are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively. The signal received at AP $m$
$\textbf{y}_{m}=\sum\nolimits_{t=1}^{T}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
In decentralized data detection, each AP performs data detection for its served UEs using a local linear received combining vector and also each AP computes its own interference-suppressing local weight based on the local signal-to-interference-plus-noise ratio (SINR). These weights are applied to the locally detected data symbols before being sent to the CPU. This differs from c-LSFD, where local data and channel estimates are transmitted to the CPU for centralized processing. The locally weighted data symbol $\hat{x}_{mt}$ for UE $t$ at AP $m$ is given by
$\hat{x}_{mt}=a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{y}_{m},$ (3)
where $\mathbf{v}_{m{t}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for UE $t$ at AP $m$, and $a_{mt}$ is the corresponding decoding weight. These weighted estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
$\hat{x}_{t}=\sum\nolimits_{m\in\mathcal{M}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{y}_{m}.$ (4)
After expanding $\textbf{y}_{m}$, the (4) can also be rewritten as
$\begin{split}\hat{x}_{t}&=\!\!\!\!\!\!\sum_{m\in\mathcal{M}_{t}}\!\!\!\!\!\!\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\hat{\textbf{g}}_{mt}x_{t}+\!\!\!\!\!\!\sum_{m\in\mathcal{M}_{t}}\!\!\!\!\!\!\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\tilde{\textbf{g}}_{mt}x_{t}\\ &+\sum_{\begin{subarray}{c}k=1\\ k\neq t\end{subarray}}^{T}\sum_{m\in\mathcal{M}_{t}}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{g}_{mk}x_{k}+\!\!\!\!\!\sum_{m\in\mathcal{M}_{t}}\!\!\!a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{n}_{m}.\end{split}$ (5)
Based on (5), the achievable uplink SE is obtained using the bounding technique and the corresponding Shannon capacity lower bound [15, 16, 17] is is given by Theorem 1.
@@
$\text{where,}\ \ \ \ \textsc{DS}_{t}=$ $\sum\nolimits_{m\in\mathcal{M}_{t}}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{t}}^{H}\hat{\textbf{g}}_{mt}\right\},$
$\textsc{UI}_{tk}=$ $\sum\nolimits_{m\in\mathcal{M}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{g}_{mk},$
$\textsc{GN}_{t}=$ $\sum\nolimits_{m\in\mathcal{M}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{n}_{m}.$
@@
First, before establishes the relationship between c-LSFD and d-LSFD, we first propose Lemma 1.
Lemma 1.
For any $q\geq 0$, define
$F(x,q)\triangleq\frac{|1+x|^{2}}{|1+x|^{2}+(q-|x|^{2})},\quad x\in\mathbb{C}.\ \ \text{Then}$
$\inf_{|x|^{2}\leq q}F(x,q)=\begin{cases}1-q,&0\leq q<1,\\[5.69054pt] 0,&q\geq 1.\end{cases}$
Moreover, for $0\leq q<1$, the minimum is attained at x=qx=-q, while for $q\geq 1$ the infimum 0 is approached as $x\to-1$.
Proof: See Appendix -B.
The following theorem establishes the relationship between c-LSFd and d-LSFD weights and quantifies the performance loss incurred by decentralization.
Theorem 2.
: The c-LSFD weights and the d-LSFD weights that maximize the SE lower bound of UE $t$ are given by $\mathbf{a}_{t}^{\mathrm{o}}=p^{u}_{t}\mathbf{J}_{t}^{-1}\mathbf{d}_{t}$ and $\mathbf{a}_{t}^{\mathrm{d}}=p^{u}_{t}\mathbf{D}_{t}^{-1}\mathbf{d}_{t}$ respectively, where,
$\mathbf{J}_{t}=\sum\nolimits_{k=1}^{T}p^{u}_{k}\mathbb{E}\{\textbf{u}_{t}\textbf{u}^{H}_{t}\}-p^{u}_{t}\textbf{d}_{t}\textbf{d}^{H}_{t}+\textbf{F}_{t}+\mathbf{\tilde{D}}_{t},$
$\text{where},$
$\textbf{d}_{t}{=}\!\Big[\!\mathbb{E}\{\!\mathbf{v}_{1{t}}^{H}\mathbf{g}_{1t}\!\},{\ldots},\!\mathbb{E}\{\!\mathbf{v}_{M{t}}^{H}\mathbf{g}_{Mt}\!\}\!\Big]^{T}\!\!\!\!,\ \!\textbf{u}_{tk}{=}\Big[\!\mathbf{v}_{1{t}}^{H}\mathbf{g}_{1k},{\ldots},\!\mathbf{v}_{M{t}}^{H}\mathbf{g}_{Mk}\!\Big]^{T}$
$\textbf{F}_{t}=\operatorname{diag}\!\Big(\Big[\mathbb{E}\{|\mathbf{v}_{1{t}}^{H}\mathbf{n}_{1}|^{2}\},\;\ldots,\;\mathbb{E}\{|\mathbf{v}_{M{t}}^{H}\mathbf{n}_{M}|^{2}\}\Big]^{T}\Big).$
Here $\mathbf{D}_{t}{=}\mathrm{diag}(\mathbf{J}_{t})$ and $\mathbf{\tilde{D}}_{t}\in\mathbb{R}^{M\times M}$ is the diagonal matrix with $\mathbf{\tilde{D}}_{t}]_{m,m}=1$, if $m\notin\mathcal{M}_{t}$ and zero otherwise. Define $\rho_{t}\triangleq\frac{\|\mathbf{a}_{t}^{\mathrm{d}}-\mathbf{a}_{t}^{\mathrm{o}}\|_{2}}{\|\mathbf{a}_{t}^{\mathrm{o}}\|_{2}}$, and the condition number of $\mathbf{J}_{t}$ as $\kappa_{t}\triangleq\frac{\lambda_{\max}(\mathbf{J}_{t})}{\lambda_{\min}(\mathbf{J}_{t})}$. Then the achievable SINR of d-LSFD satisfies
$\frac{\mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{d}})}{\mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{o}})}\;\geq\;\max\!\left\{0,\;1-\kappa_{t}\rho_{t}^{2}\right\}.$ (8)
Proof: See Appendix -B.
Remark 2: The bound in (8) shows that the performance gap between d-LSFD and c-LSFD depends on the relative weight mismatch $\rho_{t}$ and the conditioning of $\mathbf{J}_{t}$. In practice, the dominant factor is the mismatch $\rho_{t}$, which arises due to local nature of d-LSFD weights. Consequently, d-LSFD approximates the centralized solution, and when $\rho_{t}$ is small the resulting SINR loss becomes negligible.
The following proposition highlights a special case in which the d-LSFD becomes identical to c-LSFD.
Proposition 1 (Exactness of the bound).
: If the local combining scheme at each AP is either MR or ZF and each UE is assigned an orthogonal pilot sequence ($L_{p}=T$), then the interference covariance matrix $J_{t}$ becomes diagonal. Consequently, the d-LSFD weights coincide with the c-LSFD weights, i.e., $\mathbf{a}_{t}^{d}=\mathbf{a}_{t}^{o}$, and the SINR bound in Theorem 2 holds with equality.
Proof: See Appendix -C.
Remark 3: For ZF/MR combining, the approximation of d-LSFD depends on pilot contamination as non-diagonal term of $\mathbf{J}_{t}$ depends on pilot contamination. This implies that lesser the pilot contamination better the approximation.
III Adaptive Combining Analysis
In this section, we focus on G-PFZF and G-PWPFZF schemes that overcome the limitations of threshold-based UE partitioning. For these combining schemes, we derive closed-form expressions for the uplink SE and the corresponding d-LSFD weights.
@@ III Adaptive Combining Analysis
III-A Generalized Partial Full-Pilot Zero-Forcing Combining Scheme
In this work, we propose the G-PFZF combining scheme, in which each AP independently designs combining vectors at the pilot level. Since APs cannot distinguish between individual UEs sharing the same pilot, all UEs associated with the same pilot employ a common pilot-level combining vector. The pilot-level combining design is governed by a local sum spectral efficiency optimization that determines whether each pilot is treated as strong or weak, thereby explicitly controlling the allocation of spatial degrees of freedom between interference suppression and desired signal enhancement. Unlike conventional PFZF, the proposed G-PFZF does not enforce that at least one pilot must be treated as strong, thereby allowing each AP to flexibly operate in MR, PFZF, or full ZF mode depending on local channel and interference condition.
@@ III-A Generalized Partial Full-Pilot Zero-Forcing Combining Scheme
For AP $m$, let $\mathcal{S}_{m}$ and $\mathcal{W}_{m}$ denote the sets of strong and weak pilots, respectively, and let $L_{\mathcal{S}_{m}}$ be the number of strong pilots. Define the set of strong pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}{=}\{r_{m,1},\dots,r_{m,L_{\mathcal{S}_{m}}}\}$. Let $\bar{\mathbf{G}}_{m}{\in}\mathbb{C}^{A\times L_{p}}$ collect the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilots, and define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}{=}[\mathbf{e}_{r_{m,1}},{\dots},\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]$. For a given pilot $i{\in}\mathcal{S}_{m}$, let $j_{mi}{\in}\{1,2,{\dots},L_{\mathcal{S}_{m}}\}$ denote the index of pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}{\in}\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}{=}\mathbf{e}_{i}$. The pseudo-inverse of matrix for ZF structure can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$. To obtain the combining vector associated with a specific pilot, this pseudo-inverse term is multiplied by $\bm{\varepsilon}_{j_{mi}}$ and subsequently normalized. Thus, the local ZF vector for pilot $i{\in}\mathcal{S}_{m}$ at AP $m$ is given by:
@@
Following [18, 19, 20], the normalization term in (9), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix ($\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$) with DoF satisfying $A\geq L_{\mathcal{S}_{m}}+1$, is
@@
$\begin{split}&\mathbb{E}\Big\{\!\Big|(\!\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\!\Big|^{2}\!\Big\}=\!\!\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (13)
@@
By using the combining vectors in (9) and (10), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by (16),
@@
The closed-form expression of d-LSFD weights for UE $t$ at $m$ for G-PFZF combining scheme following the Theorem 2 for SINR expression in (16) is given by:
$a_{mt}=\frac{\sqrt{p^{{u}}_{t}}(b_{tt}^{m})}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}},$ (18)
where,
@@
${W}_{{m}t}=\sum\nolimits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
The derivation of the closed-form SINR and d-LSFD expressions are given in Appendix -D.
Remark 4: In conventional PFZF, UEs are classified as strong or weak using LSFCs thresholds. However, because multiple UEs may share the same pilot, UE-level binary grouping does not yield a unique or explicit mapping to the number of spatial degrees of freedom required for interference suppression. In contrast, pilot-level binary grouping establishes a explicit mapping, allowing the number of required spatial degrees of freedom to be written explicitly as $L_{\mathcal{S}_{m}}=\sum_{i=1}^{L_{p}}\delta_{mi}$, which is essential for formulating a tractable optimization problem.
Adaptive Pilot Grouping for G-PFZF Combining: This subsection details the local optimization framework that each AP employs to independently partition its set of pilots into strong and weak.
The optimization is based on a local SINR metric derived from the global SINR expression in (16). Specifically, the local SINR of UE $t$ at AP $m$ is obtained by assuming that only AP $m$ contributes to the reception of UE $t$. Moreover, by substituting $L_{\mathcal{S}_{m}}{=}\sum_{i=1}^{L_{p}}\delta_{mi}$, the dependence of the SINR on the pilot-level grouping decision at AP $m$ becomes explicit. This explicit representation is essential, as the pilot grouping variables $\{\delta_{mi}\}$ constitute the optimization variables in the proposed framework. The resulting local SINR expression is given in (19).
@@
Allocating a pilot to the strong group ($\delta_{mi}{=}1$) assigns a ZF combining vector to that pilot, thereby suppressing coherent interference from UEs sharing the same pilot as well as from UEs associated with other strong pilots. This interference suppression is achieved by consuming one spatial DoF per strong pilot, which reduces the available array gain to $(A{-}L_{\mathcal{S}_{m}})$ for all UEs served by AP $m$. Conversely, classifying a pilot as weak preserves the full array gain but provides no interference suppression for the corresponding UEs. This fundamental trade-off between array gain and interference mitigation makes the pilot grouping problem non-trivial and renders fixed, network-wide threshold-based grouping highly suboptimal. To navigate this trade-off, AP $m$ independently solves a local optimization problem, which is formulated as
@@
where $\bm{\delta}_{m}=[\delta_{m1},\delta_{m2},\cdots,\delta_{mL_{p}}]^{\intercal}$. The constraint (20c) represents the necessary condition of Wishart matrix for interference suppression. Also, intuitively, this would otherwise lead to invalid negative values for the desired signal power and coherent interference in the SINR expression.
The problem in (20) is a Mixed-Integer Non-Linear Program (MINLP), a class of problems known for its exponential computational cost. To develop a more tractable formulation, we begin by relaxing the binary constraint (20b) into a continuous one:
@@
To ensure original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the complementary constraint:
@@
Although the problem can be solved using generic non-convex optimization solvers, such approaches become computationally expensive for large networks. Therefore, to solve this problem with low computation cost, we employ gradient based proximal gradient ascent (PGA) method [21, 22]. We first reformulate the problem as: